OpenStudy (anonymous):

Show that these expressions are equivalent: (1) 4x + 3 + 7x + 8 + 9x + 2 (2) 6x + 4 + 8x + 5 + 6x + 4

OpenStudy (solomonzelman):

add like terms on each side. Do you know what like terms are, or want a similar example?

OpenStudy (solomonzelman):

I mean set: \(\large\color{black}{4x + 3 + 7x + 8 + 9x + 2=6x + 4 + 8x + 5 + 6x + 4 }\) if you want to show that 2 expressions are equivalent to each other. THEN add like terms on both sides.

OpenStudy (solomonzelman):

need help, say so please...

OpenStudy (anonymous):

Like terms?

OpenStudy (solomonzelman):

I see you need an example, let me get on out my markers;)

OpenStudy (anonymous):

Haha :)

OpenStudy (solomonzelman):

wait, can you first tell me a simple thing: what would `2x+3x` be equivalent to?

OpenStudy (solomonzelman):

(Just want to make sure, I don't go one step ahead of your current skill set)

OpenStudy (solomonzelman):

2x+3x=5x, would it be so?

OpenStudy (solomonzelman):

got disconnected, lost my work

OpenStudy (solomonzelman):

\(\Large\color{slate}{ \color{brown}{4x + 3 + 7x + 8 + 9x + 2}=\color{teal}{6x + 4 + 8x + 5 + 6x + 4 } }\) \(\Large\color{slate}{ \color{brown}{\color{blue}{4}x+\color{blue}{7}x+\color{blue}{9}x+\color{green}{3}+\color{green}{8}}+\color{green}{2}=\color{teal}{\color{red}{6}x+\color{red}{8}x+\color{red}{6}x }+\color{purple}{4}+\color{purple}{4}+\color{purple}{5} }\)

OpenStudy (solomonzelman):

this is what I would start from

OpenStudy (solomonzelman):

so you add all the: ~ x's together on each side ~ numbers together on both sides